Simplify the expression. Include absolute value bars where necessary.
4nth root t^4

Answer by tutorcecilia(2152)

The 4th root of t^4 is t because:
t*t*t*t=t^4
"t" is taken four times so the root is "t"
.
Example:
the 2nd root of 9 or 3^2 is 3 because 3*3 =9. "3" is taken two times so the root is "3".